OpenStudy (anonymous):

how come e^jπ=-1 ?????????

OpenStudy (anonymous):

depends on what you think \[e^x\] is

OpenStudy (anonymous):

if you are using \[e^{i\theta}=\cos(\theta)+i\sin(\theta)\] you get the answer by substitution

OpenStudy (anonymous):

no its fixed...as euler pointed out

OpenStudy (anonymous):

well if it ain't broke, don't fixed it

OpenStudy (amistre64):

\[e=\sum_{k=0}^{\infty}\frac{1}{k!}\]